OpenStudy (anonymous):

1. F(X)=6^X 2.F(X)=(1/7)^X 3.F(X)=4^X+^2 4.F(X)=(1/3)^X-3

OpenStudy (anonymous):

does anyone know how to graph expoential funcions

OpenStudy (anonymous):

f(x) = 6^x f(0) = 6^0 = 1 f(1) = 6^1 = 6 f(2) = 6^2 = 36 so points to plot (0,1),(1,6),(2,36) U can do more points. Same sort of thing with other functions.